Analyze the problem, answer it on a short bond paper. Follow GRESA ( Given, Required, Equation, Solution and Answer) method.

PROBLEM:

A typhoon Ompong traveled 60 km, 65° north of east and then changed its direction to 55 km, 30° south of west. What is the average velocity of the typhoon if it traveled for 24 hours?
1
Expert's answer
2020-08-24T11:56:56-0400
"\vec{s}=\vec{s_1}+\vec{s_2}\\\vec{s}=(60\cos{65},60\sin{65})-(55\cos{30},55\sin{30})\\\vec{s}=(-22.27,26.88)\ km\\\vec{v}=\frac{1}{24}(-22.27,26.88)\ km"

The average velocity of the typhoon is 1.45 kmph, west of north.